### v1

# A Unified Solution to Diverse Heterogeneities in One-shot Federated Learning

Abstract.
One-shot federated learning (FL) limits the communication between the server and clients to a single round, which largely decreases the privacy leakage risks in traditional FLs requiring multiple communications. However, we find existing one-shot FL frameworks are vulnerable to distributional heterogeneity due to their insufficient focus on data heterogeneity while concentrating predominantly on model heterogeneity. Filling this gap, we propose a unified, data-free, one-shot federated learning framework (FedHydra) that can effectively address both model and data heterogeneity. Rather than applying existing value-only learning mechanisms, a structure-value learning mechanism is proposed in FedHydra. Specifically, a new stratified learning structure is proposed to cover data heterogeneity, and the value of each item during computation reflects model heterogeneity. By this design, the data and model heterogeneity issues are simultaneously monitored from different aspects during learning. Consequently, FedHydra can effectively mitigate both issues by minimizing their inherent conflicts. We compared FedHydra with three SOTA baselines on four benchmark datasets. Experimental results show that our method outperforms the previous one-shot FL methods in both homogeneous and heterogeneous settings.
One-shot Federated Learning, Data Free, Heterogeneity Problems
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Distributed artificial intelligence
1. Introduction
Federated learning (FL) can promise enhanced privacy and data security via collaborative effort in training a global model across multiple clients without sharing private data (McMahan et al., 2017). Existing studies implicate two limitations in traditional FL, i.e., (I) it is vulnerable to potential attacks due to its reliance on multi-round communications (Rao et al., 2024; Wu et al., 2023; Zhang et al., 2019; Jere et al., 2020); (II) the performance will be largely impaired under heterogeneity scenarios (Mendieta et al., 2022; Huang et al., 2022; Karimireddy et al., 2020; Acar et al., 2020). These issues highlight the need for heterogeneity-tolerant and short-communicated FLs to ensure privacy and efficiency across diverse and sensitive applications.
One-shot FL is proposed to overcome limitation (I) by limiting the communications into one-round. By this design, one-shot FLs can minimize risks associated with data sharing and communication costs while maintaining data privacy. Meanwhile, existing one-shot FLs show potential capabilities to address limitation (II) (Guha et al., 2019; Dennis et al., 2021; Li et al., 2021b; Heinbaugh et al., 2022; Zhou et al., 2020; Zhang et al., 2022a). For example, recent methods have used auxiliary public datasets (Dennis et al., 2021; Guha et al., 2019; Li et al., 2021b; Zhou et al., 2020) or embedding client data (Heinbaugh et al., 2022) to mitigate the heterogeneity issue. Despite its potential, one-shot FL still performs poorly in highly heterogeneous scenarios, which largely prevents one-shot FL from being applied in reality (Mendieta et al., 2022; Huang et al., 2022; Qu et al., 2022; Tang et al., 2022b; Zhu et al., 2021).
To further discover the cause for low heterogeneity tolerance in existing one-shot FLs, we review the heterogeneity scenarios from data or model dimensions. Data heterogeneity can be presented as label heterogeneity i.e., Fig. 1 (a), which arises when clients have different classes (He, 2017) or size heterogeneity i.e., Fig. 1 (b), which refers to the disparity in dataset volumes across clients (He and Garcia, 2009). The label heterogeneity can skew global model performance towards more prevalent classes, while size heterogeneity leads to the problem of dominance by larger datasets. Model heterogeneity i.e., Fig. 1 (c) involves clients employing diverse model architectures (Li et al., 2020).
Diverse heterogeneity scenarios might present simultaneously but existing solutions can only address heterogeneity from one dimension. For example, auxiliary public datasets are introduced to mitigate data heterogeneity (Dennis et al., 2021; Guha et al., 2019; Li et al., 2021b; Zhou et al., 2020) but the model heterogeneity is unsolved. This will limit their application where customized models are required to meet specific needs. Recent data-free one-shot FLs, FEDCVAE-ENS and FEDCVAE-KD (Heinbaugh et al., 2022), although managing heterogeneity well, have high risks of data leakage as they directly embed the client data and upload the embedding to the server. DENSE, another data-free one-shot FL, known for its practical applicability, eliminates the need for auxiliary datasets and can accommodate model heterogeneity (Zhang et al., 2022a). However, data heterogeneity remains an unaddressed challenge that weakens its performance.
Figure 1. Illustration of heterogeneity problems (a) Label Heterogeneity, (b) Size Heterogeneity, and (c) Model Heterogeneity.
To fill this gap, we propose FedHydra, a data-free one-shot FL framework that can solve data and model heterogeneity problems. Existing studies assume all clients equally contribute to the global model in the existing studies and thus apply a value-only learning mechanisms to aggregate heterogeneous client models on the server. However, this can not handle the data heterogeneity scenarios. To include data heterogeneity during aggregation, we propose a structure-value learning mechanism where the structure should be able to simulate the impact of data heterogeneity on the global model.
As the data heterogeneity must be from clients, its impact first goes through the client and then directs to the server. To mimic this client-server pathway, we propose a new stratified structure so that implementing the value-learning mechanism within this structure can overcome data and model heterogeneity during aggregation. In addition, the data heterogeneity is monitored by the stratified structure and the model heterogeneity is monitored by the value-learning. By this design, FedHydra can effectively mitigate data and model heterogeneity issues by minimizing their inherent conflicts.
The contributions of the proposed FedHydra are as follows: 1) A structure-value learning mechanism is proposed to simultaneously monitor data and model heterogeneity. 2) A stratified structure is proposed to mimic the pathway of the impact of data heterogeneity 3) We extensively compare FedHydra across multiple benchmark datasets under both homogeneous and heterogeneous settings. The results consistently demonstrate the superior efficacy of our method compared to state-of-the-art baselines.
2. Related Works
FL (McMahan et al., 2017), introduced as a decentralized learning framework, is facing the challenge of the heterogeneity issue. Various strategies have been developed to tackle this issue, including enhancements to the FedAvg algorithm (McMahan et al., 2017) with proximal terms to control divergent updates (Li et al., 2020; Karimireddy et al., 2020; Wang et al., 2020; Acar et al., 2020; Li et al., 2021a) and the use of knowledge distillation (KD) to navigate parameter averaging challenges presented by non-IID data, either by utilizing auxiliary data (Lin et al., 2020; Sattler et al., 2021) or generative models (Zhu et al., 2021; Zhang et al., 2022b) for efficient learning across client models. Recent efforts have also aimed at optimizing client selection (Tang et al., 2022a), mitigating catastrophic forgetting to safeguard against the loss of learned information (Huang et al., 2022), and broadening the applicability of client models (Mendieta et al., 2022). Nevertheless, these strategies primarily cater to the traditional FL model, depending heavily on repeated communications, which contrasts sharply with the one-shot FL paradigm that restricts exchanges to a singular round, highlighting a gap in addressing statistical heterogeneity within this constrained communication context.
One-shot FL shows promising performance due to its minimal communication, which applies strategies like heuristic client selection for ensemble learning and KD techniques utilizing auxiliary datasets for aggregation (Guha et al., 2019). Subsequent research expanded KD applications to accommodate diverse local model structures (Li et al., 2021b). Despite achieving high accuracy, these one-shot FL methods often depend on auxiliary data, challenging their application in scenarios where data cannot be freely shared. Innovations such as dataset distillation (Zhou et al., 2020) and data augmentation (Shin et al., 2020) have been explored for secure data transfer, with Zhang et al. (Zhang et al., 2022a) proposing a data-free KD approach leveraging a generator network trained on client model ensembles. Heinbaugh et al. (Heinbaugh et al., 2022) introduced the concepts of FEDCVAE-ENS and FEDCVAE-KD to tackle statistical heterogeneity by leveraging knowledge distillation, respectively, to enhance the efficiency of model training. However, this method manipulates the local data on clients, and the uploaded decoder may cause high risks of data leakage. Nevertheless, many one-shot FL methods struggle with or have not been tested against high statistical heterogeneity, which is the gap our proposed methods aim to fill.
In data-free KD (Liu et al., 2021), DeepInversion (Yin et al., 2020) create images by optimizing RGB pixels using cross-entropy and regularization losses, enhancing the quality of the synthesis by minimizing feature statistics in Batch Normalization (BN) layers. DAFL (Chen et al., 2019) generates images using a generator with the teacher model acting as a discriminator. ZSKT (Micaelli and Storkey, 2019) generates images highlighting discrepancies between the student and teacher models to avoid overfitting in the student model. Unlike methods that select the optimal image for KD training, our strategy leverages all generated intermediate images to mitigate overfitting risks. Furthermore, although Raikward et al. (Raikwar and Mishra, 2022) introduced a KD technique that utilizes random noise as a training input, it still necessitates using real images during the training process. Besides, it requires several iterative adjustments to the BN layer statistics. In contrast, our approach enables one-shot FL without needing real images during training.
3. Methodology
This section introduces the proposed FedHydra framework and the theory and algorithms of the Model Stratification and Structure-Value Learning Mechanism phases.
3.1. FedHydra: An overview
The proposed FedHydra presents a data-free one-shot FL training scheme comprising two core components: model stratification and a structure-value learning mechanism. The latter encompasses data generation and model distillation. Within the model stratification phase, we access the classification capability of all client models across all class labels. This involves evaluating the refined weight of each client model for a particular label derived from the training loss surface of the generator, with guidance from the local model. These refined weights, serving as indicators of classification capability, are then utilized to compute the stratified aggregation (SA) for prediction logits of client models. The SA facilitates the training of both the generator and the global model.
In the structure-value learning mechanism, the data generation phase is responsible for training a generator with the SA-based guidance of client models. The synthetic data generated by this well-trained generator is subsequently employed in realizing the distillation training of the global model. Throughout the entire training process, the two phases alternate. This ensures the alignment of knowledge between synthetic and real data. In our framework, the server generates a unified global model based on client models and a single generator within a single communication round. The learning procedure is illustrated in Fig. 2, while Algorithm 1 outlines the complete training process of FedHydra.
Figure 2. Illustration of the FedHydra framework.
Algorithm 1 FedHydra Algorithm
1:Input: Client models $\{F_{k}(\cdot)\}_{k=1}^{m}$ with parameters ${\{\theta^{k}\}_{k=1}^{m}}$, the global model $F_{g}(\cdot)$ with parameter $\theta^{g}$, the Generator $G(\cdot)$ with parameter $\theta_{G}$, class labels $\{y^{j}\}_{j=1}^{c}$, the global model training epochs $T_{g}$ and learning rate $\eta_{g}$, and the generator training epochs $T_{G}$ and learning rate $\eta_{G}$.
2:Output: Well trained global model $\theta^{g}$
3:procedure FedHydra():
4:     for each client model $\theta^{k}$ in parallel do
5:         $\theta^{k}\leftarrow\textbf{LocalUpdate}(k)$
6:         Upload $\theta^{k}$ to the server
7:     end for
8:     // model stratification
9:     $\overline{U}_{r}$, $\overline{U}_{c}\leftarrow$ ModelStratification ($\{\theta^{k}\}_{k=1}^{m}$, $\theta_{G}$, $\{y^{j}\}_{j=1}^{c}$)
10:     $\triangleright$ See Algorithm 2.
11:     Initialize $\theta^{g}$ and $\theta_{G}$
12:     for each global training epoch $t_{g}$ from 1 to $T_{g}$ do
13:         Sample a batch of noises and labels $\{z_{i},y_{i}\}_{i=1}^{b}$, denoted
14:         as $\mathbf{z}=\{z_{i}\}_{i=1}^{b}$ and $\mathbf{y}=\{y_{i}\}_{i=1}^{b}$
15:         // data generation
16:         for each generator training epoch $t$ from 1 to $T_{G}$  do
17:              Generate synthetic data $\hat{\mathbf{x}}=\{\hat{x}_{i}\}_{i=1}^{b}\leftarrow G(\mathbf{z};\theta_{G}^{t})$
18:              Compute the loss $\mathcal{L}_{Gen}$ by Eq. (16)
19:              $\triangleright$ See Algorithm 3 for SA computation.
20:              $\theta_{G}^{t+1}\leftarrow\theta_{G}^{t}-\eta_{G}\nabla_{\theta_{G}^{t}}\mathcal{L}_{Gen}$
21:         end for
22:         // model distillation
23:         Compute the loss $\mathcal{L}_{Glo}$ by Eq. (19)
24:         $\theta^{g}_{t_{g}+1}\leftarrow\theta^{g}_{t_{g}}-\eta_{g}\nabla_{\theta^{g}_{t_{g}}}\mathcal{L}_{Glo}$
25:     end for
26:     return $\theta^{g}$
27:end procedure
3.2. Model Stratification
We first conduct the model stratification, aiming to evaluate the classification capability of all client models for all class labels. Specifically, we need to obtain two relative weight matrices. One weight matrix represents the guidance capability of a particular client model across all class labels, and the other represents the guidance capability for a particular class label across all client models.
Assume there are $c$ class labels indexed by $j$ and $m$ clients indexed by $k$ in the federated task. To evaluate the classification capability of one client model for a particular class $j$, we can allow a single client model $F_{k}(\theta^{k})$ to separately guide the training of the Generator $G(\theta_{G})$ for generating data with the same class label. In the $t^{th}$ iteration, we can calculate the following cross-entropy (CE) loss:
(1) $\ell_{k,j}^{t}(\hat{x}_{t}^{j},y_{t}^{j};\theta_{G})=CE(F(\hat{x}_{t}^{j};\theta^{k}),y_{t}^{j}),$
where $\hat{x}_{t}^{j}$ and $y_{t}^{j}$ denotes the generated samples and assigned labels. After the entire $T_{G}$ iterations, we can obtain the corresponding loss variance denoted as $L_{k,j}=[\ell_{k,j}^{1},..,\ell_{k,j}^{t},..,\ell_{k,j}^{T_{G}}]$. Since larger loss variance and smaller minimum loss value represent the stronger guidance to the Generator from the teacher model, the guidance capability of the client model $F_{k}(\theta^{k})$ to the class $j$ can be defined as
(2) $u_{k,j}=\frac{max(L_{k,j})-min(L_{k,j})}{min(L_{k,j})}.$
Based on this, we can obtain a basic guidance capability matrix with a size of $c\times m$ for all client models across all class labels as
(3) $U=([(u_{k,j})]_{m\times c})^{T}.$
Each row of $U$ showcases the guidance capability for a single class label across all client models and each column of $U$ represents the guidance capability of a single client model across all class labels. To realize the stratified aggregation for prediction logits of client models, we need to compare the measure values in $U$ to attain relative importance weights of guidance capability of a particular client model across all class labels as well as for a particular class label across all client models. So we first normalize each row of $U$ to $[0,1]$ by
(4) $\overline{u}_{k,j}=\frac{u_{k,j}}{\sum_{k=1}^{k=m}u_{k,j}}.$
The final normalized $U$ by row can be defined as follows:
(5) $\overline{U}_{r}=([(\overline{u}_{k,j})]_{m\times c})^{T},$
wherein $\sum_{k=1}^{k=m}\overline{u}_{k,j}=1$. Each row vector of $\overline{U}_{r}$ represents the stratified weights for prediction logits from all client models for a particular class sample.
Then, we normalize each column of $U$ to $[0,1]$ as follows:
(6) $\widehat{u}_{k,j}=\frac{u_{k,j}}{\sum_{j=1}^{j=c}u_{k,j}}.$
The final normalized $U$ by column is defined as
(7) $\overline{U}_{c}=([(\widehat{u}_{k,j})]_{m\times c})^{T},$
wherein $\sum_{j=1}^{j=c}\widehat{u}_{k,j}=1$. Each column vector of $\overline{U}_{c}$ represents the stratified weights for prediction logits of a particular client model across all class labels. The detailed procedure for conducting the model stratification is provided in Algorithm 2.
Algorithm 2 Model Stratification
1:Input: Client models ${\{F_{k}(\theta^{k})\}_{k=1}^{m}}$, the Generator $G(\theta_{G})$ and class labels $\{y^{j}\}_{j=1}^{c}$
2:Output: Stratified weight matrices $\overline{U}_{r}$ and $\overline{U}_{c}$
3:procedure ModelStratification(${\{\theta^{k}\}_{k=1}^{m}}$, $\theta_{G}$, $\{y^{j}\}_{j=1}^{c}$):
4:     for each client model $\theta^{k}$, $k$ from $1$ to $m$  do
5:         for each class label $y^{j}$, $j$ from $1$ to $c$ do
6:              Initialize the Generator parameter $\theta_{G}$
7:              Sample a batch of noise vectors and labels
8:     $\{z_{i},y_{i}^{j}\}_{i=1}^{b}$, denoted as $\mathbf{z}=\{z_{i}\}_{i=1}^{b}$ and $\mathbf{y}=\{y_{i}^{j}\}_{i=1}^{b}$
9:              // evaluate the guidance capability by Generator
10:              for each training epoch $t$ from 1 to $T_{G}$ do
11:                  $\hat{\mathbf{x}}=\{\hat{\mathbf{x}}_{i}\}_{i=1}^{b}\leftarrow$ $\textit{G}(\mathbf{z};\theta_{G}^{t})$
12:                  $\ell_{k,j}^{t}\leftarrow CE(F_{k}(\hat{\mathbf{x}};\theta^{k}),\mathbf{y})$
13:                  $L_{k,j}\leftarrow\ell_{k,j}^{t}$
14:                  $\theta_{G}^{t+1}\leftarrow\theta_{G}^{t}-\eta_{G}\nabla_{\theta_{G}^{t}}\ell_{k,j}^{t}$
15:              end for
16:              Compute $u_{k,j}$ based on $L_{k,j}$ by Eq. (2)
17:              $U\leftarrow u_{k,j}$
18:         end for
19:     end for
20:     Compute $\overline{U}_{r}$ and $\overline{U}_{c}$ by Eq. (5) and Eq. (7) based on $U$
21:     return $\overline{U}_{r}$ and $\overline{U}_{c}$
22:end procedure
3.3. Structure-Value Learning Mechanism
3.3.1. Stratified Aggregation
[…]
Based on Eq. (7), we first conduct an in-model weighted batch prediction logits $\hat{P}_{k}$ for each $P_{k}$ by
(8) $\hat{P}_{k}=P_{k}\odot(\mathbf{I}_{b\times 1}\times(\overline{U}_{c}(:,k))^{T}),$
wherein $\odot$ means the Hadamard product, $\mathbf{I}_{b\times 1}$ represents a column vector of size $b\times 1$ filled with ones, and $\overline{U}_{c}(:,k)$ denotes the $k^{th}$ column of $\overline{U}_{c}$. Then, we can obtain in-model weighted prediction logits set $P^{i}$ for the $i^{th}$ in-batch sample across all client models as below:
(9) $P^{i}=\begin{bmatrix}\hat{P}_{1}(i,:)\\ ..\\ \hat{P}_{k}(i,:)\\ ..\\ \hat{P}_{m}(i,:)\\ \end{bmatrix}_{m\times c}(i=1,2,..,b),$
wherein $\hat{P}_{k}(i,:)$ means the $i^{th}$ row of $\hat{P}_{k}$. Subsequently, based on Eq. (5), the weight set for the guidance capability of all client models on batch samples can be calculated by
(10) $V=\begin{bmatrix}\overline{U}_{r}(\textbf{y}(1),:)\\ ..\\ \overline{U}_{r}(\textbf{y}(i),:)\\ ..\\ \overline{U}_{r}(\textbf{y}(b),:)\\ \end{bmatrix}_{b\times m},$
wherein $\textbf{y}(i)$ denotes the $i^{th}$ target class label and $\overline{U}_{r}(\textbf{y}(i),:)$ means the $\textbf{y}(i)^{th}$ row of $\overline{U}_{r}$. Combining Eq. (9) and Eq. (10), the inter-model weighted batch prediction logits $P$ can be calculated by
(11) $P=\begin{bmatrix}V(1,:)\times P^{1}\\ ..\\ V(i,:)\times P^{i}\\ ..\\ V(b,:)\times P^{b}\\ \end{bmatrix}_{b\times c},$
wherein $V(i,:)$ represents the $i^{th}$ row of $V$, and $P$ is the final ensemble prediction logits output, termed Stratified Aggregation (SA). The specific computation process is summarized in Algorithm 3.
3.3.2. Data Generation
In this phase, our objective is to train a generator guided by client models to generate synthetic data with a similar distribution that closely matches the distribution of the group client data. Inspired by (Zhang et al., 2022a), we consider the generator training from three aspects.
We first compute the SA prediction logits with Algorithm 3 for a batch-generated data $\hat{\mathbf{x}}=G(\mathbf{z};\theta_{G})$ and their assigned labels $\mathbf{y}$, as follows:
(12) $P=\textbf{{SA}}(\hat{\mathbf{x}},\mathbf{y},\{\theta^{k}\}_{k=1}^{m}),$
where $P$ is our proposed SA logits. Then, to ensure the distribution similarity between synthetic data and real client data, we minimize the cross-entropy (CE) loss between $P$ and y, defined in Eq. (13), in the training objective function of the generator.
(13) $\mathcal{L}_{CE}(\hat{\mathbf{x}},\mathbf{y};\boldsymbol{\theta}_{G})=CE(P,\mathbf{y}).$
Second, Batch Normalization (BN) loss introduced in (Zhang et al., 2022a) is also employed to increase the stability of generator training, defined as follows:
(14) $\mathcal{L}_{BN}(\hat{\mathbf{x}};\boldsymbol{\theta}_{G})=\frac{1}{m}\sum_{k=1}^{m}\sum_{l}\left(\|\mu_{l}(\hat{\mathbf{x}})-\mu_{k,l}(\hat{\mathbf{x}})\|+\left\|\sigma_{l}^{2}(\hat{\mathbf{x}})-\sigma_{k,l}^{2}(\hat{\mathbf{x}})\right\|\right),$
where $\mu_{l}(\hat{\mathbf{x}})$ and $\sigma_{l}^{2}(\hat{\mathbf{x}})$ are the batch-wise mean and variance of hidden features from the $l^{th}$ BN layer of the generator, $\mu_{k,l}(\hat{\mathbf{x}})$ and $\sigma_{k,l}^{2}(\hat{\mathbf{x}})$ are the mean and variance of the $l^{th}$ BN layer of the $k^{th}$ client model.
Algorithm 3 Stratified Aggregation (SA)
1:Input: Client models ${\{F_{k}(\theta^{k})\}_{k=1}^{m}}$, two weight matrices $\overline{U}_{r}$ and $\overline{U}_{c}$ from model stratification, and synthetic batch data $\mathbf{(\hat{x},y)}=\{x_{i},y_{i}\}_{i=1}^{b}$
2:Output: weighted batch prediction logits matrix $P$
3:procedure $\textbf{{SA}}({\{\theta^{k}\}_{k=1}^{m}},\overline{U}_{r},\overline{U}_{c},\mathbf{(\hat{x},y)})$
4:     for each client model $\theta^{k}$, $k$ from $1$ to $m$ do
5:         $P_{k}\leftarrow F_{k}(\textbf{x};\theta^{k})$
6:         Compute in-model weighted prediction logits $\hat{P}_{k}$ based
7:   on $P_{k}$ and $\overline{U}_{c}$ by Eq. (8)
8:     end for
9:     Compute inter-model weighted prediction logits $P$ based on
10:  $\{\hat{P}_{k}\}_{k=1}^{m}$ and $\overline{U}_{r}$ by Eq. (9), (10), and (11)
11:     return $P$
12:end procedure
Third, to increase the diversity of generated data, an adversarial distillation (AD) loss is considered to maximize the discrepancy between SA logits and prediction logits of the global model, defined as follows:
(15) $\mathcal{L}_{AD}(\hat{\mathbf{x}},\mathbf{y};\boldsymbol{\theta}_{G})=-KL\left(P,F_{g}(\hat{\mathbf{x}};\boldsymbol{\theta}^{g})\right),$
where $F_{g}(\cdot)$ and $\theta^{g}$ denote the global model and its parameter. Therefore, the final generator training loss is defined by combining the above three loss terms, as follows:
(16) $\mathcal{L}_{Gen}(\hat{\mathbf{x}},\mathbf{y};\boldsymbol{\theta}_{G})=\mathcal{L}_{CE}(\hat{\mathbf{x}},\mathbf{y};\boldsymbol{\theta}_{G})+\lambda_{1}\mathcal{L}_{BN}(\hat{\mathbf{x}};\boldsymbol{\theta}_{G})+\lambda_{2}\mathcal{L}_{AD}(\hat{\mathbf{x}},\mathbf{y};\boldsymbol{\theta}_{G}),$
where $\lambda_{1}$ and $\lambda_{2}$ are adjustable parameters.
3.3.3. Model Distillation
Based on the synthetic data produced in the data generation phase, we design a new model distillation scheme to train our global model with knowledge distillation techniques.
We consider the two loss terms in the training objective function of the global model. First, we use the KL loss between the SA logits and prediction logits of the global model to distill the knowledge from client models to the global model, defined as
(17) $\mathcal{L}_{KL}(\hat{\mathbf{x}},\mathbf{y};\boldsymbol{\theta}^{g})=KL\left(P,F_{g}(\hat{\mathbf{x}};\boldsymbol{\theta}^{g})\right).$
To achieve efficient and stable model distillation, ensuring consistent prediction labels between the ensemble client models and the global model is crucial, especially considering the potential deviation between synthetic and real data. So, we introduce the following CE loss, defined as:
(18) $\mathcal{L}_{\widehat{CE}}(\hat{\mathbf{x}},\mathbf{y};\boldsymbol{\theta}^{g})=CE\left(F_{g}(\hat{\mathbf{x}};\boldsymbol{\theta}^{g}),H[P]\right),$
where $H[\cdot]$ represents to obtain the hard label of prediction logits. Therefore, by combining the above two loss terms, the final training loss function for realizing model distillation is defined as follows:
(19) $\mathcal{L}_{Glo}(\hat{\mathbf{x}},\mathbf{y};\boldsymbol{\theta}^{g})=\mathcal{L}_{KL}(\hat{\mathbf{x}},\mathbf{y};\boldsymbol{\theta}^{g})+\beta\mathcal{L}_{\widehat{CE}}(\hat{\mathbf{x}},\mathbf{y};\boldsymbol{\theta}^{g}),$
where $\beta$ is an adjustable parameter.
4. Experiments
4.1. Setup
4.1.1. Datasets
Our evaluation experiments utilize four datasets: MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), SVHN (Netzer et al., 2011), and CIFAR10 (Krizhevsky et al., 2009). MNIST comprises 60,000 training images and 10,000 testing images of handwritten digits, each with a size of 28x28 pixels. FashionMNIST provides a more challenging fashion item classification dataset, with the same quantities of image and label and size specifications as MNIST. SVHN consists of over 600,000 color images of house numbers extracted from Google Street View, suitable for digit recognition tasks. Lastly, CIFAR10 comprises 60,000 32x32 color images across 10 object classes, with 50,000 images for training and 10,000 for testing.
4.1.2. Client Data Distribution
We adopt the Dirichlet distribution method ($Dir(\alpha)$) in (Yurochkin et al., 2019) to simulate diverse client data distributions in practical FL scenarios, as shown in Fig. 3. Adjusting the parameter $\alpha$ allows us to modify the level of data heterogeneity: a lower $\alpha$ results in more pronounced heterogeneity in the data distribution, while a higher $\alpha$ results in a homogeneous data distribution. We establish $\alpha=0.5$ by default to initiate this process.
Figure 3. Visualization of heterogeneous distributions of client data characterized by four diverse $\alpha$ values, with MNIST used as an illustration. The horizontal axis denotes the client ID, while the vertical axis represents the class label. Each circle’s size corresponds to the number of data samples linked to a specific class, and circles of identical color denote the same class label.
4.1.3. Baselines
Since the proposed method aims to solve the heterogeneity problem in one-shot FL, we omit the comparison with some approaches, like FedBE (Chen and Chao, 2020), and FedGen (Zhu et al., 2021), which need to download auxiliary models or datasets. In addition, methods such as FedProx (Li et al., 2020), FedNova (Wang et al., 2020), and Scaffold (Karimireddy et al., 2020) are also not considered because regularization-based methods are not designed for one-shot FL. Furthermore, we are not comparing with FEDCVAE (Heinbaugh et al., 2022), which directly embeds the data on the client side and uploads it to the server without model training on the client. Therefore, We choose three State-of-the-art (SOTA) baselines: FedAvg (McMahan et al., 2017), FedDF (Lin et al., 2020), and DENSE (Zhang et al., 2022a).
4.1.4. Configurations
In our FL setup, we set the client number $m=5$. For the client model training, we employ the SGD optimizer with learning rate $\eta=0.01$, while setting the local epochs and batch size as $B=128$ and $E=200$, respectively. On the server side, we utilize the Adam optimizer with learning rate $\eta_{G}=0.001$ to train the generator. This generator, characterized by adjustable parameters $\lambda_{1}=1.0,\lambda_{2}=0.5$, undergoes training for $T_{G}=30$ rounds to generate batch examples. For the distillation training of the global model, we set the training epochs to $T_{g}=200$ and the learning rate to $\eta_{g}=0.01$, with SGD serving as the optimizer.
4.2. Evaluation
We report the key experimental results in this section, and more experiment results are reported in the appendix within the submitted Supplementary Material.
4.2.1. Performance Overview
Table 1 presents the testing performance on FedHydra and comparative methods over four datasets with different $\alpha$ of client data distribution, which varies from $0.5$ to $0.01$ wherein $\alpha=0.01$ indicates an extreme data heterogeneity situation. The results show that our FedHydra achieves the highest accuracy across all the data heterogeneity settings with all the datasets. Specifically, all the methods perform well on the MNIST dataset because of less data heterogeneity when $\alpha=0.5$. Our FedHydra outperforms the baseline FedAvg by $3.99$%, FedDF by $6.85$%, and DENSE by $7.22$%, respectively. Furthermore, When the data heterogeneity increases, we observe that FedAvg decreases dramatically from $81.89$% to $69.89$% when $\alpha=0.3$, then further down to $25.96$% when $\alpha=0.1$, the accuracy only achieved $10.10$% when $\alpha=0.01$ under an extremely data heterogeneity situation. Other methods like FedDF and DENSE show similar trends when the data heterogeneity increases. However, our FedHydra shows robustness under data heterogeneity situations, which outperforms the second-best method, DENSE, by $7.22$% when $\alpha=0.3$, $23.24$% when $\alpha=0.1$, and $27.66$% when $\alpha=0.01$. This is because FedAvg averages the uploaded parameters from the clients, which causes less robustness when data heterogeneity happens. FedDF requires prior data with the average ensemble logits from the uploaded models to process the model distillation, similar to the reason for DENSE. The averaged ensemble logits could not guide the generator to generate the synthetic data by specific labels when data heterogeneity increases.
The result patterns are similar on other datasets, including FashionMNIST, SVHN, and CIFAR10. Our FedHydra outperforms the second-highest by $1.44$%, $1.58$%, $12.24$% and $4.37$%, and achieved $77.20$%, $76.39$%, $50.84$% and $46.02$% from $\alpha=0.5$ to $0.01$ on FashionMNIST, respectively. On SVHN, FedHydra outperforms the second-highest by $2.19$%, $1.96$%, $7.73$% and $9.01$%, and achieved $82.39$%, $82.96$%, $62.08$% and $30.95$% from $\alpha=0.5$ to $0.01$. In addition, FedHydra also outperforms the second-highest by $4.67$%, $6.09$%, $5.56$% and $7.94$%, and achieved $63.15$%, $57.66$%, $48.12$% and $40.63$% from $\alpha=0.5$ to $0.01$ on CIFAR10, respectively. Therefore, the results prove that the proposed FedHydra significantly outperforms the SOTA baseline methods, further illustrating the superior performance of FedHydra in a data heterogeneity environment.
Table 1. Top-1 test accuracy achieved by FedAvg, FedDF, DENSE, and FedHydra over various datasets with varying $\alpha$.
Dataset MNIST FashionMNIST SVHN CIFAR10
Method $\alpha$=0.5 $\alpha$=0.3 $\alpha$=0.1 $\alpha$=0.01 $\alpha$=0.5 $\alpha$=0.3 $\alpha$=0.1 $\alpha$=0.01 $\alpha$=0.5 $\alpha$=0.3 $\alpha$=0.1 $\alpha$=0.01 $\alpha$=0.5 $\alpha$=0.3 $\alpha$=0.1 $\alpha$=0.01
FedAvg 81.89 69.89 25.96 10.10 75.76 74.81 30.71 33.71 59.94 64.64 31.46 21.94 38.26 25.54 23.87 14.96
FedDF 78.23 73.59 31.52 10.10 59.38 65.61 32.11 40.36 80.20 81.00 54.35 7.59 54.81 50.82 33.06 19.91
DENSE 77.86 76.78 36.73 11.17 71.14 72.79 38.24 41.65 79.44 78.81 53.69 7.75 58.48 51.57 42.56 32.69
FedHydra (ours) 85.08 85.16 59.97 38.83 77.20 76.39 50.48 46.02 82.39 82.96 62.08 30.95 63.15 57.66 48.12 40.63
Further experiments show the impact of accuracy versus model distillation epochs over four datasets. The corresponding test accuracy curves in each distillation epoch are provided in Fig. 4 and 5, respectively. In Fig. 4. FedHydra significantly outperforms the SOTA baseline methods after $25$ epochs, especially when $\alpha\leq 0.3$. In Fig. 5, FedHydra demonstrates dominant capabilities on the CIFAR10 dataset, significantly outperforming the SOTA baseline methods from $25$ epochs to $200$ with $\alpha=\{0.5,0.3,0.1,0.01\}$.
Figure 4. Test accuracy versus distillation epoch over FashionMNIST with $\alpha=\{0.5,0.3,0.1,0.01\}$.
Figure 5. Test accuracy versus distillation epoch over CIFAR10 with $\alpha=\{0.5,0.3,0.1,0.01\}$.
4.2.2. Impact of Stratified Aggregation
DENSE employs a simple averaging ensemble to aggregate the output logits from client models. This means that the averaging ensemble equally treats the classification capability of all client models on diverse class samples. However, different client models may exhibit varying classification capabilities across different class examples, especially in FL settings with data heterogeneity. Hence, our FedHydra initially evaluates the stratified weights of classification capabilities among various client models across diverse class labels. These weights are then applied to the prediction logits of client models to achieve stratified aggregation, which guides the structure-value learning mechanism.
To validate the effectiveness of stratified aggregation in heterogeneous federated data distributions, we designed a more extreme data distribution scenario with only two disjoint class samples per client, termed 2c/c. Also, taking MNIST with 2c/c distribution as an example, Fig. 6 (a) illustrates the data distribution across five clients in the 2c/c case. Following this federated client data distribution, each updated client model exhibits strong classification capability solely on examples belonging to its exclusively possessed two class labels. This alignment is well-reflected in the stratified weight distributions, evaluated by model stratification, of the classification capability of the five client models across the ten class labels, as depicted in Fig. 6 (b). For instance, important classification weights of client 1 for class labels 0 and 1 are $0.9594$% and $0.9797$%, respectively, while weights for other class labels are nearly zero. Hence, our model stratification accurately assesses the classification capabilities of client models.
Table 2 shows the top-1 test accuracy on four datasets with 2c/c distribution. Under the extreme data distribution case, FedHydra still shows the dominant performance over the SOTA baseline methods by overtaking the second-highest method by $20.31$%, $2.33$%, $8.58$% and $7.73$%, which achieves $40.81$%, $39.06$%, $50.61$% and $39.09$% on MNIST, FashionMNIST, SVHN, and CIFAR10, respectively. FedDF and DENSE are relatively lower than FedAvg on simple datasets under 2c/c distribution such as MNIST and FashionMNIST because averaging ensemble logits will fail to provide valuable guidance for data generation and model distillation. However, due to the model parameter sensitivity to complicated classification tasks, FedAvg shows poor performance on complex datasets, such as SVHN and CIFAR10. The results prove that our FedHydra is robust under 2c/c distribution. Perceptive test accuracy variations are plotted in Fig. 7. The results show that our FedHydra outperforms all the other methods over four datasets under 2c/c distribution.
Furthermore, we compare the test accuracy of the averaging ensemble (AE) from DENSE and our stratified aggregation (SA) over four datasets, as shown in Fig. 8. The results indicate that our SA is much higher than AE in every numerical $\alpha$, including 2c/c. AE can reach $20.00$% less than SA when $\alpha=0.5$ and $0.3$, where the degree of data heterogeneity is not that high. However, when $\alpha=0.1$, AE decreases greatly from $81.75$% to $37.41$% on MNIST, for example. The gap is accompanied by the increased degree of data heterogeneity, which reaches the largest gap when $\alpha=0.01$, the degree of data heterogeneity $\alpha=0.01$ is higher than 2c/c because $\alpha=0.01$ represents both label and size heterogeneity. In contrast, 2c/c size are equal.
Table 2. Top-1 test accuracy achieved by FedAvg, FedDF, DENSE, and FedHydra over various datasets with 2c/c distribution.
Dataset MNIST FashionMNIST SVHN CIFAR10
FedAvg 20.50 36.73 26.68 9.22
FedDF 13.38 15.33 42.03 17.43
DENSE 15.58 22.31 41.38 31.36
FedHydra (ours) 40.81 39.06 50.61 39.09
Figure 6. Visualization of stratified aggregation. (a) 2c/c data distribution across five clients using MNIST as an example. (b) The weight distribution of evaluated classification capability for five client models on ten class labels.
Figure 7. Test accuracy versus distillation epoch over diverse datasets with 2c/c distribution.
Figure 8. Test accuracy achieved by averaging ensemble (AE) and our stratified aggregation (SA) over four datasets with varying client data distributions.
4.2.3. Impact of Stratified Aggregation Guided Model Distillation
We demonstrate the impact of our SA-guided model distillation by comparing it with FedAvg, FedDF, and DENSE. Taking CIFAR10 with $\alpha=0.1$ as an example, the results are plotted in Fig. 9. Fig. 9 (a) shows the results of the global model and local models across different local training epochs $E=\{40,80,120,...,400\}$. The global model achieves the best performance around $24.00$% when $E=40$. This is because of the mismatch in optimization goals when dealing with data heterogeneity, resulting in weight divergence. However, Fig. 9 (b) represents the FedAvg, FedDF, DENSE, and FedHydra results ranging $E$ from 00 to $400$. FedHydra outperforms DENSE and each client model, while FedAvg performance is under client models. The results prove that SA-guided model distillation is robust under data heterogeneity while SA-guided model distillation harms the distillation performance.
Figure 9. (a) Test accuracy of FedAvg and client models across local training epochs $E={20,40,60,...,400}$. (b) Local training accuracy of different clients. The dotted lines represent the top-1 test accuracy achieved by FedAvg, FedDF, DESNE, and our FedHydra.
Table 3. Top-1 test accuracy achieved by FedDF, DENSE, and FedHydra over CIFAR10 with personalized client models.
Data Distribution Personalized Client Model Server (ResNet18)
GoogLeNet ResNet18 CNN1 CNN2 LeNet FedDF DENSE FedHydra
$\alpha$=0.01 26.68 28.78 26.71 19.28 10.00 22.25 18.67 30.51
$\alpha$=0.1 37.36 36.63 35.60 30.41 28.77 27.39 26.43 30.90
$\alpha$=0.3 46.41 55.49 45.78 47.98 33.47 46.96 44.54 48.80
$\alpha$=0.5 49.33 61.08 47.65 54.14 48.25 56.27 51.29 59.23
4.2.4. Performance in FL Model Heterogeneity
Our proposed FedHydra can simultaneously adapt to data heterogeneity and model heterogeneity. We design five different client models, including GoogleNet, ResNet18, two simple CNNs, and LeNet, for local training on CIFAR10 by varying $\alpha$ values in $\{0.01,0.1,0.3,0.5\}$ to evaluate the performance under both model and data heterogeneity. We choose ResNet18 as the global model on the server side. FedAvg does not support model heterogeneity. Thus, we compare our FedHydra with FedDF and DENSE. Table 3 summarises the results of different models over CIFAR10 under both data and model heterogeneity settings. FedHydra outperforms the other methods on the server, which achieves $59.23$% overtaking FedDF by $2.96$%. However, by the decrease of $\alpha$, both FedDF and DENSE drop dramatically from $56.27$% to $22.25$% and $51.29$% to $18.67$% when $\alpha=0.01$. FedHydra still promises robustness in model and data heterogeneity, achieving $30.51$%.
5. Conclusion
We propose FedHydra, a unified data-free one-shot federated learning framework designed to effectively address both data and model heterogeneity. FedHydra differentiates itself by employing a structure-value learning mechanism instead of the conventional value-only approach. It incorporates a newly devised stratified learning structure that accounts for data heterogeneity, and each learning item included in this structure reflects model heterogeneity. This innovative design allows for the concurrent management of data and model heterogeneity, enabling FedHydra to efficiently alleviate challenges posed by both heterogeneity. Comparative evaluations of FedHydra against SOTA baselines across four benchmark datasets demonstrate its superior performance in both homogeneous and heterogeneous settings, outstripping previous one-shot FL methodologies.
Appendix A More Performance Overview Results on MNIST and SVHN
Figure 10. Test accuracy versus distillation epoch over MNIST with $\alpha=\{0.5,0.3,0.1,0.01\}$.
Fig. 10 shows the curves of accuracy versus distillation epoch over MNIST with $\alpha=\{0.5,0.3,0.1,0.01\}$, the results show that FedHydra overtakes other baseline methods after round $15$ epochs, and the proposed FedHydra fluctuates but still can reach the peak at $60$ epochs around $35.00$% when $\alpha=0.01$, while others keep the performances as $10$%.
Figure 11. Test accuracy versus distillation epoch over SVHN with $\alpha=\{0.5,0.3,0.1,0.01\}$.
On the SVHN dataset, FedHydra outperforms other baseline methods after $50$ epochs as shown in Fig. 11. Additionally, FedHydra shows a dithering upward trend and peaks around $60$ epochs when $\alpha=0.01$.
Appendix B Extended Analysis
B.1. Impact of Client Numbers
Furthermore, we also test the impact of client numbers on SVHN and CIFAR10 by varying the numbers of clients $K=\{5,10,20,50,100\}$, as shown in Table 4. All the methods decrease when the client number increases. However, FedHydra has less impact than other methods, which achieves $60.01$% when $K=100$ and outperforms the second-best FedDF with only $50.37$% on the SVHN dataset. The same trends are shown on the CIFAR-10 dataset. FedHydra achieves $47.40$% when $K=100$ and outperforms the second-best DENSE with only $36.25$%. Thanks to the proposed SA algorithm, FedHydra minimally affects the increase in clients.
Table 4. Top-1 test accuracy achieved by FedAvg, FedDF, DENSE, and FedHydra over SVHN and CIFAR10 with different client numbers $K=\{5,10,20,50,100\}$.
Client SVHN CIFAR10
$K$ FedAvg FedDF DENSE FedHydra FedAvg FedDF DENSE FedHydra
5 59.94 80.20 79.44 82.39 38.26 54.81 58.48 63.15
10 43.45 70.20 69.51 77.13 41.71 50.17 55.43 61.58
20 47.87 69.36 69.10 76.54 33.47 38.70 42.50 59.64
50 36.77 68.30 67.74 74.89 31.75 37.85 39.09 52.66
100 23.69 50.37 48.87 60.01 27.23 33.90 36.25 47.40
B.2. Performance Under Multiple Global Training Rounds
We extend our experiments to multiple global rounds. Specifically, we fix the local epoch $E=50$ and $\alpha=0.1$ to evaluate the performance of diverse methods on SVHN and CIFAR10 with varying client distributions. Table 5 summarizes the relevant evaluation results. The results in Table 5 show that FedHydra outperforms the other methods with different global training rounds from $T=\{1,2,3,4,5\}$. In SVHN, the performance of FedHydra steadily increases and reaches a peak when $T=3$, which outperforms DENSE by $1.86$%. In CIFAR-10, FedHydra still performs better than other methods, which keeps steady after $T=2$ and overtake the second-best FedAvg by $9.2$%.
Table 5. Top-1 test accuracy achieved by FedAvg, FedDF, DENSE, and FedHydra over SVHN and CIFAR10 with $\alpha=0.1$ under multiple global rounds.
SVHN CIFAR10
Global
round
FedAvg FedDF DENSE FedHydra FedAvg FedDF DENSE FedHydra
$T$=1 28.31 48.13 48.22 53.27 22.72 21.33 29.60 37.95
$T$=2 34.55 56.62 54.67 59.98 37.32 28.63 32.39 46.52
$T$=3 44.60 65.61 68.44 70.30 35.95 30.36 36.42 48.75
$T$=4 57.05 63.62 66.83 68.58 43.86 30.71 38.85 50.17
$T$=5 61.59 66.66 64.57 70.39 41.99 28.82 37.27 49.34
B.3. Visualization of Synthetic Data
To assess how synthetic data compares to actual training data, we provide the image comparison between the original training data and synthetic data on FashionMNIST and CIFAR-10 in Fig. 12. The figure demonstrates that the first and third rows display the original data from FashionMNIST and CIFAR10, respectively. In contrast, the second and last rows showcase the synthetic data produced by models trained on the FashionMNIST and CIFAR-10 datasets. Notably, the synthetic data differ significantly from the original data, effectively minimizing the risk of exposing sensitive client information. Despite the marked visual differences, the synthetic data contribute to our method outperforming other baseline approaches in performance. This highlights that synthetic data should ideally be easily distinguishable from real data while supporting superior model training outcomes.
Figure 12. Visualization of synthetic data generated by the generator about FashionMNIST and CIFAR10.
Appendix C More Experiment Results on SVHN in Heterogeneous FL
Table 6 provides the test results on SVHN in heterogeneous FL. The results show that FedHydra achieves $71.29$% when $\alpha=0.5$, outperforming the second-best FedDF with $69.75$%. However, when data heterogeneity increases, FedHydra still has $54.33$% while FedDF decreases to $43.82$%.
Table 6. Top-1 test accuracy achieved by FedDF, DENSE, and FedHydra over SVHN with personalized client models.
Data Distribution Personalized Client Model Server (ResNet18)
GoogLeNet ResNet18 CNN1 CNN2 LeNet FedDF DENSE FedHydra
$\alpha$=0.01 24.17 7.59 29.81 34.76 19.52 43.82 42.39 54.33
$\alpha$=0.1 42.28 14.84 39.64 34.80 31.17 37.43 36.00 47.45
$\alpha$=0.3 66.78 65.62 38.22 58.73 64.19 77.22 75.75 77.52
$\alpha$=0.5 72.54 70.72 39.74 69.24 64.56 69.75 64.67 71.29
Table 7 summarizes the extra test results on SVHN and CIFAR10 with 2c/c in heterogeneous FL. We observe that the performance of all the client models is relatively low under the 2c/c setting, while FedDF and DENSE can only achieve $24.72$% and $20.42$% on SVHN, respectively. However, FedHydra still can reach $31.22$, which overpasses the FedDF $6.5$%. The same results are shown on CIFAR-10, FedHydra reaches $38.57$ which overpasses the FedDF $7.52$%
Table 7. Top-1 test accuracy achieved by FedDF, DENSE, and FedHydra over SVHN and CIFAR10 with 2c/c in heterogeneous OFL.
Dataset with 2c/c Personalized Client Model Server (ResNet18)
GoogLeNet ResNet18 CNN1 CNN2 LeNet FedDF DENSE FedHydra
SVHN 25.78 26.30 18.48 15.13 11.85 24.72 20.42 31.22
CIFAR10 19.43 18.10 18.69 19.62 19.00 31.05 26.18 38.57
Appendix D More Experiment Results Under Multiple Global Rounds Across Diverse $\alpha$ Values
Table 8, 9, D, and LABEL:table:mr-2cc provide the relevant test results with $\alpha=\{0.5,0.3,0.01\}$ and 2c/c under five global rounds.
Table 8. Top-1 test accuracy achieved by FedAvg, FedDF, DENSE, and FedHydra over SVHN and CIFAR10 with $\alpha=0.5$ under multiple global rounds.
Dataset SVHN CIFAR10
Global
round
FedAvg FedDF DENSE FedHydra FedAvg FedDF DENSE FedHydra
$T$=1 58.88 80.55 79.50 84.97 38.26 48.14 50.23 53.48
$T$=2 71.36 84.22 82.88 85.37 41.71 49.26 52.58 57.09
$T$=3 85.73 85.76 84.21 86.41 45.42 48.71 55.39 62.75
$T$=4 84.12 85.04 83.53 86.46 49.81 48.86 53.81 61.83
$T$=5 83.84 85.20 84.60 86.81 51.75 49.70 53.49 64.71
Table 8 shows that the proposed FedHydra outperforms and is more stabilized than all the other methods from different global rounds when $\alpha=0.5$. The performance of FedHydra keeps above $80$% on SVHN and outperforms the best second method $2$% to $3$%. The gap is much higher on CIFAR-10, where FedHydra can outperform more than $10$%, compared with the second-best DENSE on the $5_{th}$ global round. Table 9 presents that FedHydra consistently outperforms the other methods on both datasets, demonstrating a marked advantage, particularly on the more complex CIFAR10 dataset. For instance, it peaks at 86.72% on SVHN and 58.43% on CIFAR-10, significantly higher than its competitors. The $\alpha$ is set to $0.01$ on Table D, FedHydra consistently outperforms the other algorithms across all rounds and both datasets, demonstrating exceptional adaptability and robustness under high data heterogeneity. Specifically, FedHydra’s accuracy peaks at 39.58% on SVHN and 39.76% on CIFAR10, significantly surpassing the next-best results from other models.
Table 9. Top-1 test accuracy achieved by FedAvg, FedDF, DENSE, and FedHydra over SVHN and CIFAR10 with $\alpha=0.3$ under multiple global rounds.
Dataset SVHN CIFAR10
Global
round
FedAvg FedDF DENSE FedHydra FedAvg FedDF DENSE FedHydra
$T$=1 61.17 77.99 76.76 79.14 33.72 37.56 41.54 50.01
$T$=2 73.50 84.21 82.68 84.83 37.56 38.94 46.44 57.00
$T$=3 84.29 84.49 84.81 85.56 36.04 40.73 47.10 58.18
$T$=4 84.13 85.04 83.82 86.02 41.73 41.10 46.20 58.43
$T$=5 84.96 85.89 84.62 86.72 44.26 41.03 48.55 55.97
Table LABEL:table:mr-2cc assesses the performance under a 2c/c setting across five global training rounds. FedHydra consistently outshines the other algorithms in both datasets, showing its superior adaptability and robustness, with peak accuracies reaching 65.76% on SVHN and 38.76% on CIFAR10 by the fifth round. FedDF generally performs well on SVHN, albeit less consistently on CIFAR10, while DENSE exhibits mixed results across rounds and datasets. FedAvg consistently underperforms, highlighting its limitations in more complex or skewed data environments. Overall, FedHydra emerges as the most effective method, particularly adept at handling challenges presented by the 2c/c setting, suggesting strong capabilities in managing diverse and complex data distributions in federated learning scenarios. These results underscore FedHydra’s potential effectiveness in federated learning scenarios where managing diverse data distributions is critical, marking it as a superior choice in handling complex data heterogeneity challenges in federated environments.
Table 10. Top-1 test accuracy achieved by FedAvg, FedDF, DENSE, and FedHydra over SVHN and CIFAR10 with $\alpha=0.01$ under multiple global rounds.
Dataset